A lab contains 24 ounces of iodine solution of strength 15%. how much pure iodine would you mix with it to get an iodine solution of strength 25%?
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of pure iodine(that would be an 100% solution)
question states***
.15*24 + 1.00x = .25(24oz + x )
solve for x
.75x = .10*24oz
x = 3.2oz is the amount of oure iodine needed
checking our answer
.15*24 + 1.00x = .25(24oz + x )
3.6 + 3.2 = 6.8